# Collapse of Irrelevant Representations (CIR) Ensures Robust and Non-Disruptive LLM Unlearning

### v1

@@ 5 Method comparisons
We compare the methods on a task of unlearning knowledge useful for bio-terrorism and cyber-warfare. We use the Weapons of Mass Destruction Proxy (WMDP) benchmark (Li et al., 2024). We chose a high-quality subset of 144 biological and 203 cyber questions. [^4] We randomly split these into development (20%) and holdout (80%) sets. All results shown are on the holdout set (112 bio and 165 cyber questions). Similarly to Deeb & Roger (2024), for each question we generated three simple sentences and use them all as the forget set. See Appendix B for question filtering criteria and generation details.
As retain sets, we use the FineFineWeb corpus (M-A-P et al., 2024) – the biology subset for WMDP-Bio and the computer_science_and_technology subset for WMDP-Cyber.
@@
We compare CIR to two popular unlearning methods. Gradient Difference (Liu et al., 2022) which maximizes the cross-entropy loss on the forget set while minimizing the loss on the retain set, and Circuit Breakers (Zou et al., 2024) which we described in Section 4.
@@
We use the Llama-3.1-8B model (Meta, 2024). We control for the disruption of general performance (as measured by the loss on WikiText (Merity et al., 2016)), by terminating the unlearning when disruption crosses a certain threshold. Afterwards, we perform a 100 epoch fine-tuning attack, on facts different than the evaluated ones but from the same distribution. For this we use the same WMDP split as Deeb & Roger (2024), with unlearning on 100% of the data, then relearning on 80% and evaluating the accuracy on the remaining 20%. Following Sondej et al. (2025), to stabilize training we always normalize the norm of unlearning updates to some fixed value. This value effectively acts as the unlearning rate. We describe hyperparameter tuning and compute requirements in Appendix C.
We found CIR easier to tune, because it has a wider range of valid hyperparameters. In Circuit Breakers and Gradient Difference, unlearning and retaining seem to push against each other, and small changes of hyperparameters can tip the balance. As we see on Figure 4, the balance can even tip during one run, with unlearning gains getting abruptly reverted. Again, this is likely caused by these methods unlearning general representations which are also present in the retain set, so training on the retain set pushes in the opposite direction.
@@
We terminate CIR when the WikiText loss crosses 100.1% of its initial value. When we tuned the baselines using the same 100.1% threshold, none achieved meaningful accuracy decreases (the highest was only 1 percentage point), so to better assess their performance we gave them a 30x handicap (termination at 103%) and tuned them again.
Results
To measure post-attack accuracy, we smooth the relearning curve to remove noise and then take its maximum value, since some attacks were longer than optimal. Despite 30× less performance disruption, for WMDP-Bio CIR reduces post-attack accuracy 80× more than the best baseline (Figure 1(b)), and for WMDP-Cyber 30× more (Figure 4). Unlike prior methods, CIR is selective enough to be used even without retain training, although with worse performance – 51% post-attack accuracy on bio and 41% on cyber (when allowing 0.1% disruption like before).
Gradient Difference performs poorly mainly because its retain set training struggles to prevent disruption on WikiText, or often even on the evaluation split of the retain set.
On Figure 4 we show what happens if we let CIR disrupt more too (up to 1%). Surprisingly, the gains in post-attack accuracy are disproportionately low, which supports our findings from Section 3.1 about disruptive unlearning being unhelpful.  [^5] Another explanation would be that unlearning has already stopped since we have hit random accuracy level. But the probability of generating the harmful answer (not shown here) keeps decreasing, meaning that unlearning still proceeds – it is just unrobust.
@@ 6 Limitations
WMDP imprefections
We suspect that some amount of unrobustness is caused by certain WMDP questions being easy to guess without knowing the answer. For example in WMDP-Cyber, among questions where the attack increases accuracy, the correct answer is the longest option 52% of the time, compared to 14% for the rest. More information in Appendix D.
@@
In our work we targeted facts present in the WMDP dataset. But for full bio and cyber safety we will need to unlearn orders of magnitude more. Now, a significant limiting factor becomes a lack of high-quality unlearning data. Creating such datasets will require a ton of work from bio and cyber experts, and releasing them publicly would pose a security risk, so both creation and usage of such datasets will need to be coordinated for example by AI Safety Institutes.
@@
Note that the assumption that common representations are irrelevant, works well when unlearning knowledge – the relevant representations are fact-specific, and so quite rare. But if we hope to unlearn tendencies (like power-seeking, deceptiveness, etc.), then the harmful representations are often quite common across training texts. So there, choosing which representations to collapse will need to be more elaborate than simply doing PCA. We leave it for future work to explore this.
@@ Appendix A Unrelated Facts Disruption and Language Transfer
On Figure 8 we show additional examples, and we can see that disruption happens also if we ask the questions differently, without using the word ”is”. We can also see that more distant facts are disrupted less, around 8%. Still, 8% is quite severe – for example if we were to unlearn 100 facts, and their disruption was independent, then some random unrelated fact would suffer $\approx 800\%$ disruption (recall loss increase). (But it would likely be less, because their disruption would not be independent.)
We also see that there is some language transfer, but it is significant (about 50%) only for languages with similar words (”ist”, ”es”). In contrast, for Russian and Portuguese the transfer is quite weak, which would necessitate doing the unlearning in other languages too. This is consistent with a finding by Thibodeau (2022) that unlearning (in his case, the ROME technique (Meng et al., 2023)) is quite specific to the exact tokens used (for example unlearning facts about ”cheese”, does not transfer to ”fromage”).
@@ Appendix B Unlearning corpus creation
We started off with a subset of WMDP created by Deeb & Roger (2024), where they filtered out skill-based questions and duplicates (WMDP-Deduped). Then, for faithful answer recall evaluations, we wanted to create a dataset where the answer can be cleanly separated from the non-harmful context, but we found that many answers were convoluted and long, containing mostly benign tokens. So we kept only the questions with answers shorter than 60 characters. We also excluded ”none of the above” and ”all of the above” answers, because they lead to awkward generated forget corpus.
@@
Generation prompt asks for simplicity and not adding unnecessary text. In Table 1 we see it indeed produces simpler sentences than in the best corpus from Deeb & Roger (2024) who used a similar generation approach. (But for most questions the improvement was smaller than in the table.) We saw that this avoiding of unnecessary text greatly reduces disruption from unlearning. We invite future designers of unlearning corpora to also include only essentially harmful text.
@@
Prior corpus from Deeb & Roger (2024) $\bullet$ In the synthesis of RNA virus genomes, obtaining virions solely from sequence information poses the potential risk of re-creating and engineering dangerous pathogens. This concern highlights the importance of careful consideration and oversight in the manipulation and study of viral genetic material.

### v2

@@ 5 Experimental setup
We evaluate unlearning methods on bio-terrorism and cyber-warfare knowledge using the Weapons of Mass Destruction Proxy (WMDP) benchmark (li_wmdp_2024). We selected a high-quality subset of 144 biological and 203 cyber questions. [^5] We randomly split these into development (20%) and holdout (80%) sets. All results are reported on the holdout set (112 bio and 165 cyber questions). Following deeb_unlearning_2024, we generate three simple sentences per question and use them as the forget set. Filtering and generation details are in Appendix B. As retain sets we use the FineFineWeb corpus (finefineweb2024): the biology subset for WMDP-Bio and the computer_science_and_technology subset for WMDP-Cyber.
@@
We compare CIR to two popular unlearning methods: Gradient Difference (liu2022continuallearningprivateunlearning), which maximizes cross-entropy on the forget set while minimizing loss on the retain set; and Circuit Breakers (zou_improving_2024), described in Section 4.
@@
We use the Llama-3.1-8B model (grattafiori2024llama). We control for disruption of general performance (measured by the loss on WikiText (merity2016pointer)) by terminating the unlearning when disruption crosses a certain threshold. After unlearning, we perform a 100 epoch fine-tuning attack on facts different than the evaluated ones but from the same distribution. We follow the WMDP split from deeb_unlearning_2024: unlearning on 100% of the data, relearning on 80%, and evaluation on the remaining 20%. Following sondej_robust_2025, we stabilize training by normalizing the norm of unlearning updates to a fixed value. This value effectively acts as the unlearning rate. Hyperparameter tuning and compute requirements are detailed in Appendix C.
@@
We terminate CIR when the WikiText loss crosses 100.1% of its initial value. When we tune the baselines using the same 100.1% threshold, none achieves meaningful accuracy decreases (the highest was only 1 percentage point), so to better assess their performance, we give them a 30x handicap (termination at 103%) and retune their hyperparameters.
6 Results
CIR is easier to tune
We found that CIR admits a wide range of valid hyperparameters. By contrast, in Circuit Breakers and Gradient Difference unlearning and retaining seem to push against each other, and small changes of hyperparameters can tip the balance. As Figure 5 shows, the balance can even flip during a single run, with unlearning gains abruptly reverting. Again, this fragility likely arises because those methods remove general representations that also appear in the retain set, so training on the retain set updates the model in the opposite direction to unlearning.
CIR outperform baselines on both robustness and non-disruption
To measure post-attack accuracy, we smooth each relearning curve to remove noise and report its peak value, since some attacks run longer than optimal. Despite 30× less performance disruption, for WMDP-Bio CIR reduces post-attack accuracy 80× more than the best baseline (Figure 1(b)), and for WMDP-Cyber 30× more (Figure 5). Unlike prior methods, CIR is selective enough to be used even without retain training, although with worse performance – 51% post-attack accuracy on bio and 41% on cyber (with the same 0.1% disruption budget as before).
Gradient Difference performs poorly mainly because its retain set training struggles to prevent disruption on WikiText and often even on the evaluation split of the retain set.
Disruptive unlearning is not robust
Figure 5 shows what happens if we let CIR disrupt more (up to 1%). Surprisingly, the gains in post-attack accuracy are disproportionately low, which supports our findings from Section 3.1 that disruptive unlearning is unhelpful. [^6] An alternative explanation is that unlearning has plateaued after hitting random accuracy; however, the probability of generating the harmful answer (not shown here) continues to decrease, indicating unlearning is still proceeding, it is just not robust.
7 Conclusion
We identified why current unlearning methods fail: they disrupt general representations shared between harmful and benign capabilities, which can be easily reversed with fine-tuning. Our Collapse of Irrelevant Representations (CIR) technique addresses this fundamental issue by precisely targeting only the representations specific to the unlearned facts. On WMDP benchmarks, CIR achieves over 30× stronger unlearning robustness than prior methods, while disrupting performance 30× less, proving that representational selectivity is essential for unlearning.
@@ 8 Limitations
WMDP imperfections
We suspect that some unrobustness is caused by certain WMDP questions being easy to guess without knowing the answer. For example, in WMDP-Cyber, among questions where the attack increases accuracy, the correct answer is the longest option 52% of the time, compared to 14% for the rest. More information in Appendix D. Fixing this issue may push post-attack accuracies even closer to the random level.
@@
In our study, we target facts present in the WMDP dataset. Scaling to full bio and cyber safety will require unlearning orders of magnitude more facts. A bottleneck to this, is the lack of high-quality unlearning data, with existing bio and cyber unlearning corpora (li_wmdp_2024) containing mostly benign text. Creating better datasets will require a ton of work from bio and cyber experts, and releasing them publicly would pose a security risk, so both creation and usage of such datasets will need careful coordination by specialized bodies.
@@
Note that the assumption that common representations are irrelevant works well when unlearning knowledge, as the relevant representations are fact-specific, and therefore relatively rare. But if we hope to unlearn tendencies (such as power-seeking, deceptiveness, etc.), then the harmful representations are often quite common across training texts. So choosing which representations to collapse will need to be more elaborate than simply doing PCA. We leave it for future work to explore.
@@ Acknowledgments
(a) Fine-grained grid search for the optimal number of projections. Uses CIR + Circuit Breakers loss on layers 6-15, with only 0.2% allowed disruption.
(b) Same as (a), but a wider search range, and 0.5% allowed disruption.
(c) Search for the optimal layers for the intervention, with 3 different algorithms. 0.5% allowed disruption.
Figure 6: CIR hyperparameter searches.
In all experiments we report WMDP-Cyber accuracy at temperature=1, after a fine-tuning attack. All the attacks have converged. For cleaner comparisons, no retain training was used. Note that 1 projected component means just projecting the mean and no actual PCs (which is efficient but performs poorly).
@@ Appendix A Unrelated Facts Disruption and Language Transfer
On Figure 7 we show additional examples, and we can see that disruption happens also if we ask the questions differently, without using the word ”is”. We can also see that more distant facts are disrupted less, around 8%.
We also see that there is some language transfer, but it is significant (about 50%) only for languages with similar words (”ist”, ”es”). In contrast, for Russian and Portuguese the transfer is quite weak, which would necessitate doing the unlearning in other languages too. This is consistent with a finding by jacques_thibodeau_but_2022 that unlearning (in his case, the ROME technique (meng_locating_2023)) is quite specific to the exact tokens used (for example unlearning facts about ”cheese”, does not transfer to ”fromage”).
@@ Appendix B Unlearning corpus creation
We started off with a subset of WMDP created by deeb_unlearning_2024, where they filtered out skill-based questions and duplicates (WMDP-Deduped). Then, for faithful answer recall evaluations, we wanted to create a dataset where the answer can be cleanly separated from the non-harmful context, but we found that many answers were convoluted and long, containing mostly benign tokens. So we kept only the questions with answers shorter than 60 characters. We also excluded ”none of the above” and ”all of the above” answers, because they lead to awkward generated forget corpus.
@@
Generation prompt asks for simplicity and not adding unnecessary text. In Table 1 we see it indeed produces simpler sentences than in the best corpus from deeb_unlearning_2024 who used a similar generation approach. (But for most questions the improvement was smaller than in the table.) We saw that this avoiding of unnecessary text greatly reduces disruption from unlearning. We invite future designers of unlearning corpora to also include only essentially harmful text.
@@
Prior corpus from deeb_unlearning_2024 $\bullet$ In the synthesis of RNA virus genomes, obtaining virions solely from sequence information poses the potential risk of re-creating and engineering dangerous pathogens. This concern highlights the importance of careful consideration and oversight in the manipulation and study of viral genetic material.
@@
Figure 8: Comparison of three ways of breaking representations.
In our method we minimize the dot product of current and initial activations, clipped at 0 to avoid the dot product becoming negative. Secondly, we tried simply minimizing the norm of the current activations. Lastly, we tried minimizing the cosine similarity between current and initial activations, also clipped at 0 – this was used in the original circuit breakers paper (zou_improving_2024).
(We used CIR, with Llama-3.1-8B and measured activation norm at layer 6.)
@@ Appendix D Guessability of WMDP questions
Appendix E Details of the disruption thresholds experiment
On Figure 2, we show 50 unlearning runs using the CIR technique. For diversity, we used many variations of CIR, using different unlearning and retaining losses (described in Section 4), different hyperparameters (learning rates, numbers of projected components), and sometimes even skipped retain training. We used Llama-3.1-8B and the WMDP-Cyber dataset. We made sure that each of the attacks has converged.
Appendix F Use of Large Language Models
In accordance with ICLR 2026 disclosure requirements, we acknowledge that large language models were used to polish writing and assist with code autocompletion during the preparation of this work. All research contributions and conclusions remain entirely the work of the authors.